\begin{thebibliography}{FGLH{\etalchar{+}}20}

\bibitem[ACJ03]{aroca2003power}
Fuensanta Aroca, Josè Cano, and Fran{\c{c}}oise Jung.
\newblock Power series solutions for non-linear pde's.
\newblock In {\em Proceedings of the 2003 international symposium on Symbolic
  and algebraic computation}, pages 15--22, 2003.

\bibitem[AGT16]{AGT}
Fuensanta Aroca, Cristhian Garay, and Zeinab Toghani.
\newblock The fundamental theorem of tropical differential algebraic geometry.
\newblock {\em Pacific Journal of Mathematics}, 283(2):257--270, 2016.

\bibitem[BFNS21]{boulier2021relationship}
Fran{\c{c}}ois Boulier, Sebastian Falkensteiner, Marc~Paul Noordman, and
  Omar~Leon Sanchez.
\newblock On the relationship between differential algebra and tropical
  differential algebraic geometry.
\newblock In {\em Computer Algebra in Scientific Computing: 23rd International
  Workshop, CASC 2021, Sochi, Russia, September 13--17, 2021, Proceedings 23},
  pages 62--77. Springer, 2021.

\bibitem[BH19]{boulier2019ritt}
Fran{\c c}ois Boulier and Mercedes Haiech.
\newblock {The Ritt-Raudenbush Theorem and Tropical Differential Geometry}.
\newblock working paper or preprint, 2019.

\bibitem[DL84]{denef1984power}
Jan Denef and Leonard Lipshitz.
\newblock Power series solutions of algebraic differential equations.
\newblock {\em Mathematische annalen}, 267:213--238, 1984.

\bibitem[Edg09]{edgar2009transseries}
Gerald~A. Edgar.
\newblock {Transseries for Beginners}.
\newblock {\em Real Analysis Exchange}, 35(2):253 -- 310, 2009.

\bibitem[FGLH{\etalchar{+}}20]{falkfund}
Sebastian Falkensteiner, Cristhian Garay-L{\'o}pez, Mercedes Haiech, Marc~Paul
  Noordman, Zeinab Toghani, and Fran{\c{c}}ois Boulier.
\newblock The fundamental theorem of tropical partial differential algebraic
  geometry.
\newblock In {\em Proceedings of the 45th International Symposium on Symbolic
  and Algebraic Computation}, pages 178--185, 2020.

\bibitem[FMSW23]{falkensteiner2023algebro}
Sebastian Falkensteiner, Johann Mitteramskogler, J~Sendra, and Franz Winkler.
\newblock The algebro-geometric method: Solving algebraic differential
  equations by parametrizations.
\newblock {\em Bulletin of the American Mathematical Society}, 60(1):85--122,
  2023.

\bibitem[GG16]{eqtrop}
Jeffrey Giansiracusa and Noah Giansiracusa.
\newblock Equations of tropical varieties.
\newblock {\em Duke Math. J.}, 165(18):3379--3433, 2016.

\bibitem[GM21]{gianmereta}
Jeffrey Giansiracusa and Stefano Mereta.
\newblock A general framework for tropical differential equations.
\newblock {\em arXiv preprint arXiv:2111.03925}, 2021.

\bibitem[Gri17]{grig}
Dima Grigoriev.
\newblock Tropical differential equations.
\newblock {\em Advances in Applied Mathematics}, 82:120--128, 2017.

\bibitem[Kol73]{kolchin1973differential}
Ellis~Robert Kolchin.
\newblock {\em Differential algebra \& algebraic groups}.
\newblock Academic press, 1973.

\bibitem[LH14]{lange2014counting}
Markus Lange-Hegermann.
\newblock {\em Counting solutions of differential equations}.
\newblock PhD thesis, Aachen, Techn. Hochsch., Diss., 2014.

\bibitem[Mer22]{tesi}
Stefano Mereta.
\newblock A general framework for tropical differential equations.
\newblock PhD thesis, Swansea University and Université Grenoble Alpes, 2022.

\bibitem[Mer23]{merfund}
Stefano Mereta.
\newblock The fundamental theorem of tropical differential algebra over
  nontrivially valued fields and the radius of convergence of nonarchimedean
  differential equations.
\newblock {\em arXiv preprint arXiv:2303.12124}, 2023.

\bibitem[MR18]{tropideal}
Diane Maclagan and Felipe Rinc{\'o}n.
\newblock Tropical ideals.
\newblock {\em Compositio Mathematica}, 154(3):640--670, 2018.

\bibitem[MS10]{moosa2010jet}
Rahim Moosa and Thomas Scanlon.
\newblock Jet and prolongation spaces.
\newblock {\em Journal of the Institute of Mathematics of Jussieu},
  9(2):391--430, 2010.

\bibitem[MS21]{macsturm}
Diane Maclagan and Bernd Sturmfels.
\newblock {\em Introduction to tropical geometry}, volume 161.
\newblock American Mathematical Society, 2021.

\bibitem[Rob14]{robertz2014formal}
Daniel Robertz.
\newblock {\em Formal algorithmic elimination for PDEs}, volume 2121.
\newblock Springer, 2014.

\bibitem[Sei10]{seiler2010involution}
Werner~M. Seiler.
\newblock {\em Involution}, volume~24.
\newblock Springer, 2010.

\bibitem[vdH01]{van2001complex}
Joris van~der Hoeven.
\newblock {\em Complex transseries solutions to algebraic differential
  equations}.
\newblock Universit{\'e} de Paris-Sud. D{\'e}partement de Math{\'e}matique,
  2001.

\bibitem[VdH06]{van2006transseries}
Joris Van~der Hoeven.
\newblock {\em Transseries and real differential algebra}, volume 1888.
\newblock Springer, 2006.

\end{thebibliography}

\documentclass[article]{amsart}

\usepackage[vmargin=3cm, hmargin=3cm]{geometry}
\usepackage[foot]{amsaddr}

\usepackage{multicol}
\usepackage{afterpage}
\usepackage{makecell}
\usepackage{enumitem}
\usepackage{hyperref}
\usepackage{mathrsfs}  
\usepackage{latexsym}
\usepackage{enumitem}
\usepackage{bbm} 
\usepackage{amsmath, amsfonts, amssymb, amsthm, tensor}
\usepackage[all]{xy}
\usepackage{verbatim}
\usepackage{tikz-cd}
\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage{multicol}
\usepackage{hhline}
\usepackage{color}
\usepackage{adjustbox}

\usepackage{mathtools}

\def\co{\colon\thinspace}

\usepackage{lineno}

\numberwithin{equation}{section}

\newtheorem{teorema}{Theorem}[section]  
\newtheorem{lem}[teorema]{Lemma} 
\newtheorem{prop}[teorema]{Proposition}

\theoremstyle{definition}
\newtheorem{defin}[teorema]{Definition}

\theoremstyle{remark} 
\newtheorem{ex}[teorema]{Example}
\newtheorem{oss}[teorema]{Remark}

\newcommand{\C}{\mathbb{C}}
\newcommand{\R}{\mathbb{R}}
\newcommand{\N}{\mathbb{N}}
\newcommand{\Z}{\mathbb{Z}}
\newcommand{\Q}{\mathbb{Q}}
\newcommand{\T}{\mathbb{T}}

\newcommand{\trop}{\textup{trop}}

\newcommand{\B}{\mathbb{B}}

\newcommand{\pk}{K[\![t]\!]}

\newcommand{\pt}{\T[\![t]\!]}
\newcommand{\pb}{\B[\![t]\!]}

\newcommand{\diff}[2]{#1\{x_1, \ldots , x_{#2}\}}

\newcommand{\Sol}{\textup{Sol}}

\newcommand{\supp}{\mathrm{supp}}
\newcommand{\ord}{\mathrm{ord}}

\newcommand{\im}{\mathrm{im}}

\title{The fundamental theorem of tropical differential algebra for formal Puiseux series}

\author{Sebastian Falkensteiner}
\email{Sebastian.Falkensteiner@mis.mpg.de} 
\address{Max Planck Institute for Mathematics in the Sciences, Inselstraße 22, 04103 Leipzig, Germany}

\author{Stefano Mereta}
\email{Stefano.Mereta@mis.mpg.de}

\date{\today}

\subjclass[2010]{Primary 14A20;
	Secondary 12H99, 13N99, 14T90, 14T99}
\keywords{tropical geometry; algebraic differential equations; fundamental theorem; Puiseux series}

\begin{document}
\begin{abstract}
The fundamental theorem of tropical differential algebra has been established for formal power series solutions of systems of algebraic differential equations. It has been shown that the direct extension to formal Puiseux series solutions fails. In this paper, we overcome this issue by transforming the given differential system and such a generalization is presented. 
Moreover, we explain why such transformations do not work for generalizing the fundamental theorem to transseries solutions, but show that one inclusion still holds for this case even without using any transformation.
\end{abstract}
\maketitle

\section{Introduction}

Even though the first appearances of tropical methods in the area of optimization can be traced back to the 1970's, it is with the beginning of the new millennium that these techniques have started to be applied more and more often, and fruitfully, to geometry, eventually developing into an independent and rich new branch of mathematics, with connections spreading as far as computational biology and machine learning. The fundamental notion linking the classical world and the tropical one is that of tropicalization: by means of a valuation, this process moves problems of algebro-geometric nature to tropical ones, the latter having an intrinsic combinatorial and polyhedral nature. A comprehensive text to approach the subject is~\cite{macsturm}.

More recently, following the revival of interest around algebraic methods for differential equations, and differential algebra in general, see e.g.~\cite{robertz2014formal,lange2014counting,falkensteiner2023algebro}, a tropical approach to the study of solutions to set of ODEs have been introduced in~\cite{grig}. In this work the author gives a definition of tropical differential equations and other objects needed to state a fundamental theorem (similar to the fundamental theorem of tropical geometry, see for example \cite[Theorem 3.2.3]{macsturm}) in this context. This theorem was successfully proven in \cite{AGT} and subsequently extended to the case of partial differential equations in \cite{falkfund}, see~\cite{boulier2021relationship} for a more algorithmic description. Drawing inspiration from recent works where tropical geometry is developed as geometry over the tropical semiring, such as \cite{eqtrop, tropideal}, a more general framework encompassing that of \cite{grig} have been proposed in \cite{gianmereta}. This allows for a tropical theory of differential equations to be developed also in the case of nontrivially valued fields of coefficients and in this context a fundamental theorem for tropical differential algebra, extending that of \cite{AGT}, have been proven in \cite{merfund}. We will state our results in this language.

The fundamental theorem of tropical differential equations states equality between the (weighted) support of formal power series solutions of a differential ideal and the tropical solutions of the corresponding tropicalized differential system over a valued uncountable algebraically closed field of characteristic zero $K$, see Theorem~\ref{thm:ft_powerseries}. 
The direct extension to more general formal series solutions such as formal Laurent series and formal Puiseux series fails as it is presented in~\cite[Section 7]{falkensteiner2023algebro}. 
In the present work, we use transformations of the differential ideal and its solution set to overcome this issue.

Transformations of differential equations and its solution sets are well studied in the literature. Of particular interest have been diffeomorphisms and homeomorphisms of dynamical systems and their solutions, the main subject of differential topology. 
For non-autonomous and algebraic differential equations, and their solutions, less is known. In these cases, transformations are usually defined simultaneously on the coordinates and the functions; for rational solutions see~\cite[Section 4]{falkensteiner2023algebro} and references therein. 
In this paper, we are loosely following the approach from~\cite{falkensteiner2023algebro} and use rational transformations. 
More precisely, we will study rational maps $\sigma(\bold{t},\bold{x})=(\bold{s}(\bold{t}),\bold{z}(\bold{t},\bold{x}))$ on the vector of independent variables $\bold{t}=(t_1,\ldots,t_m)$ and differential indeterminates $\bold{x}=(x_1,\ldots,x_n)$. In order to maintain the differential structure, we will have to perform the transformations on the derivatives of the $\bold{x}$ accordingly. This will give us a new differential system in $\bold{t}$ and $\bold{x}$. For applying the transformation in the fundamental theorem, we additionally require that
\begin{enumerate}
    \item $\bold{z}$ maps $(\bold{t},K[\![\bold{t}]\!]^n)$ injectively to $K[\![\bold{t}]\!]^n$;
    \item there has to be a tropical analogue of $\sigma$, i.e., there must be $\Sigma(\bold{t},\bold{x})=(\bold{S}(\bold{t}),\bold{Z}(\bold{t},\bold{x}))$ acting on support sets $\bold{x}$ such that $\bold{Z}(\bold{t},\trop(\varphi))=\trop(\bold{z}(\bold{t},\varphi))$ for $\varphi \in K[\![\bold{t}]\!]^n$;
    \item since we consider formal power series expanded around the origin, we require that $\psi(0)=\bold{z}(\bold{t},\varphi(\bold{t}))|_{\bold{t}=0}$ is defined and therefore assume that $0 \in \im(\bold{s})$ such that w.l.o.g. we set $\bold{s}(0)=0$; and 
    \item since $\bold{z}(\bold{t},\bold{c})$ has to be defined for every constant $\bold{c} \in K^n$, we assume that $\bold{z}(\bold{t},\bold{x})$ is polynomial in $\bold{x}$.
\end{enumerate}
For applying such transformations to univariate formal Laurent series and formal Puiseux series, we use particular choices for $\sigma$, namely the shift $\varphi \mapsto t^n \cdot \varphi$ and the power map $t \mapsto t^n$, respectively. In the multivariate case, this transformation is a bit more subtle. 
The actual computation for the value $n$ in the shift and the power map can be found for concrete differential equations for instance by the Newton polygon method for differential equations, see e.g.~\cite{aroca2003power} and references therein. Here we do not further study this finding and note that there is no general algorithm known for finding minimal values $n$.

In Section~\ref{sec-generalizationPuiseux}, we first recall the classical formulations of the fundamental theorem for tropical differential algebra. We then show the extension to formal Laurent series solutions and formal Puiseux series solutions by manipulating the given differential ideal in an appropriate way (see Theorem~\ref{thm:ft_PuiseuxseriesOrdinary} for the ordinary case and Theorem~\ref{thm:ft_PuiseuxseriesPartial} for the multivariate case). 
In Section~\ref{sec-generalizationLogarithms}, we show that such transformations $\sigma$ cannot be used for mapping transseries to formal power series. Thus, we can not generalize the fundamental theorem to transseries solutions in this way. Moreover, we show that one inclusion of the fundamental theorem still holds in the case of transseries without any transformations.
We conclude the paper by some applications of the used transformations in the setting of approximation theory.
 
\section{How to generalize the fundamental theorem to formal Puiseux series?}\label{sec-generalizationPuiseux}
The fundamental theorem of tropical (partial) differential equations is formulated and proven for (multi-variate) formal power series solutions of systems of algebraic differential equations with formal power series coefficients over an uncountable, algebraically closed fields of characteristic zero $K$. 
In the ordinary case, where these series are univariate, the coefficients of the given differential system may be generalized to formal Laurent series. 
Essentially, this can be achieved by multiplying with a monomial of lowest order occuring as a coefficient in the system. 
Note that we may assume that the given system is a prime differential ideal and finitely represented by a basis (see Ritt-Raudenbush basis theorem) and thus, this order is indeed finite. 
Since the obtained system has again formal power series coefficients, and their solution set is the same, the fundamental theorem can be applied then. 
More generally, formal Puiseux series coefficients of the given differential system can be transformed back to formal power series with a power map in the independent variable.

In this document, we instead give a generalization of the solution set to formal Puiseux series. The applied transformations, however, change the differential system and also the solution set such that we can apply the fundamental theorem for the case of formal power series solutions. 
Since this transformation is one-to-one, we can recover the original system with its formal Puiseux series solutions. 
We present a formulation of the fundamental theorem by using inverse limits.

As we want to use the language of semirings to state our definitions and results, here we make a quick recap about semirings and valuations. A \emph{semiring} $(S, \oplus, \odot)$ is an algebraic structure satisfying all the axioms to be a ring, but the existence of additive inverses. A semiring is said to be \emph{idempotent} if for every $a \in S$ we have that $a \oplus a = a$. 
\begin{ex}
\begin{enumerate}
\item The set $\B = \{0,\infty\}$ equipped with operations $\oplus := \min$ and $\odot := +$ is an idempotent semiring, called the idempotent semiring of Boolean numbers;
\item The set $\T = \R \cup \{\infty\}$ equipped with operations $\oplus := \min$ and $\odot: = +$ is an idempotent semiring, called the idempotent semiring of tropical numbers. It contains $\B$ as a subsemiring;
\item For every $n \ge 1$ let $\T_n = \R^n \cup \{\infty\}$ with $\oplus$ given by lexicographic minimum and $\odot := +$. It is an idempotent semiring, called the idempotent semiring of rank $n$ tropical numbers.
\end{enumerate}
\end{ex}
A sum of elements in an idempotent semiring is said to \emph{tropically vanish} if by deleting any of the summands its result does not change.
We adopt the following generalised definition of valuation, as introduced and used in \cite{eqtrop, gianmereta, merfund}: a \emph{valuation} on a ring $R$ is a map $v\co R \to S$ to an idempotent semiring $S$ satisfying the following conditions for all $a,b \in \T_n$:
\begin{multicols}{2}
\begin{enumerate}
\item	$v(0)  = 0_S$,
\item	$v(1) = v(-1) = 1_S$,
\item	$v(ab) = v(a) \odot v(b)$,
\item	$v(a+b) \oplus v(a) \oplus v(b)$ tropically vanishes. 
\end{enumerate}
\end{multicols}
When $S=\T_n$ for some $n$, we will say that the valuation $v$ is of \emph{rank n}. Let us note that condition (4) is equivalent to saying that the minimum of the terms $v(a+b), v(a), v(b)$ is achieved at least twice.

Let $K$ be an uncountable algebraically closed field of characteristic zero equipped with a valuation $v_K \co K \to \T$. Let $R_{m,n} = K[\![t_1,\ldots,t_m]\!]\{x_1,\ldots,x_n\}$ denote the ring of differential polynomials in $x_1,\ldots,x_n$ with formal power series coefficients. 
Moreover, letting $\bold{t}:=(t_1,\ldots,t_m)$, we write $\Sol_A$ for the set of solutions in a differential algebra $A$ over $K[\![\bold t]\!]$. 
In the following, we will use multi-index notations and operations such as multiplication, raising powers etc. are performed component-wise.

For simplicity of treatment, here we will introduce the objects and definitions appearing in the statement of the fundamental theorem of tropical differential equations only in the case of ODEs (a treatment of the case of PDEs can be found in~\cite{falkfund} when $v_K$ is the trivial valuation and in~\cite[Chapter 8]{tesi} in full generality). 
In this case, the valuation $v_K$ on $K$ induces a rank 2 valuation $v\co \pk \to \T_2$ defined as
\[
A_{n_0}t^{n_0} + \dots \mapsto (n_0, v_K(A_{n_0})) \in \T_2.
\]
When $v_K$ is the trivial valuation the image of this map is contained in a subsemiring of $\T_2$ isomorphic to $\T$, thus we recover the $t$-adic valuation used in~\cite{grig} and~\cite{AGT}.

Given a differential polynomial $f \in R_{1,n}$ of order less or equal $r \in \Z_{\ge 0}$, let us write it as $f(x) = \sum_{\lambda \in \Lambda} A_\lambda x^\lambda$ where $\Lambda$ is a finite set of matrices in $\text{Mat}_{r \times n}(\Z_{\ge 0})$, $A_\lambda$ is an element of $\pk$ for every $\lambda$ and $x^\lambda$ is the differential monomial defined as $ \prod_{i,j} ( x_i^{(j)} ) ^{\lambda_{i,j}}$. 
The tropicalization $\trop_v(f)$ of $f$ with respect to $v$ is defined as the polynomial with coefficients in $\T_2$ obtained by applying $v$ coefficient-wise to $f$:
\[
\trop_v(f) := \sum_{\lambda \in \Lambda} v(A_\lambda) x^\lambda= \sum_{\lambda \in \Lambda} (\alpha_{\lambda} \text{, } \beta_{\lambda}) x^\lambda
\] 
for the same $\Lambda$ and for $(\alpha_{\lambda} \text{, } \beta_{\lambda}) \in \R^2$ for every $\lambda \in \Lambda$. This polynomial lives in an algebra of polynomials over $\T_2$ in countably many variables $x_i^{(j)}$'s. Even though we could not talk rigorously about elements of this algebra as tropical differential polynomials over $\T_2$ (in fact this matter is quite subtle, see \cite[Section 3]{gianmereta}), we will nevertheless refer to them as tropical differential polynomials, as this will not give rise to any ambiguity. 
Given a differential ideal $G \subseteq R_{1,n}$ its tropicalization is the set $\trop_v(G) := \{\trop_v(f) \mid f \in G\}$.

Given a power series $A \in \pk$, its tropicalization is performed by applying $v_K$ coefficientwise, obtaining an element of $\pt$. We will denote this map as $\widetilde v \co \pk \rightarrow \pt $.
Lastly let $\Phi \co \pt \rightarrow \T_2$ be defined as the homomorphism of semirings
\[
B_{n_0}t^{n_0} + \dots \mapsto (n_0, B_{n_0}) \in \T_2.
\]
\begin{oss}
The valuation $v$, the homomorphism $\Phi$ and the map $\widetilde v$ fit into a commutative triangle as follows:
\begin{equation} \label{diagram:diff-enh}
\begin{tikzcd}
& \pt \arrow[d, "\Phi"]  \\
\pk \arrow[ur, "\widetilde{v}"] \arrow[r, swap, "v"] & \T_2 .
\end{tikzcd}
\end{equation}
This is an instance of \emph{differential enhancement} of the valuation $v$ (see \cite[Section 4]{gianmereta}). Furthermore notice that the maps in the diagram above can we defined in the same way considering Laurent or Puiseaux series instead of power series, giving rise to an analogous commutative diagram as above. In this case, though, the map $\Phi$ fails to be \emph{reduced}: this makes the fundamental theorem of tropical differential equations fail in this case, for similar reasons as explained in Example \ref{ex:fail-fund-theorem}.
\end{oss}
We finally introduce the notion of solution to a tropical differential equation:
\begin{defin}\label{definition:tropical-solution}
Given a differential polynomial $f \in R_{1,n}$, a vector of tropical power series $B = (B_1, \dots, B_n) \in \pt^n$ is a \emph{solution} to the tropicalization of $f$ if the expression
\[
\trop_v(f)(B) := \bigoplus_{\lambda \in \Lambda} v(A_\lambda) \bigodot_{i,j}  ( \Phi(d^j(B_i)) ) ^{\odot{\lambda_{i,j}}} \in \T_2
\]
tropically vanishes. I.e.\ if when plugging in $ \Phi(d^j(B_i))$ for $x_i^{(j)}$, the minimum is achieved at least twice in $\T_2$.
\end{defin}

As here we will not deal with the fully general case as introduced in \cite[Section 4]{gianmereta}, and deal only with the case of the homomorphism $\Phi$, we will use the notation $\Sol_{\pt}(\trop_v(f))$ for the set of solutions to the tropicalization of the differential equation $f$ in $\pt^n$, as introduced in Definition \ref{definition:tropical-solution} above.
We have the following results:
\begin{teorema}[\cite{AGT,falkfund}]\label{thm:ft_powerseries}
Let $K$ be an uncountable, algebraically closed field of characteristic zero, and let $v_K$ be the trivial valuation. Let $G$ be a differential ideal in the ring $R_{m,n}$. Then 
\[ \trop_{\widetilde v}(\Sol_{K[\![\bold{t}]\!]}(G)) = \Sol_{\T[\![\bold{t}]\!]}(\trop_v(G)). \]
\end{teorema}
In the case of ordinary differential equations the following generalisation holds:
\begin{teorema}[\cite{merfund}] \label{thm:ft_powerseries_with_coeffs}
Let $K$ be an uncountable, algebraically closed field of characteristic zero, equipped with a valuation $v_K \co K \rightarrow \T$. Let $G$ be a differential ideal in the ring $R_{1,n}$. Then 
\[ \trop_{\widetilde v}(\Sol_{K[\![\bold{t}]\!]}(G)) = \Sol_{\T[\![\bold{t}]\!]}(\trop_v(G)). \]
\end{teorema}
In the second case, we suspect that an analogous theorem for PDEs should hold. To simplify the treatment in the following, we will focus on the trivially valued case from here on, this way we will be able to treat both the case of ODEs and PDEs simultaneously.

\subsection{Ordinary case}
Throughout the section, let $G$ be a differential ideal in the ring $R_{1,n}$. 

\subsubsection*{Laurent series}
Let us look for formal Laurent series solutions of $G$ instead of formal power series solutions. A univariate formal Laurent series $\varphi = \sum_{i \ge 0} c_it^{i-l}$ with $c_i \in K$, $l \in \Z$, can be transformed to the formal power series $t^{l} \cdot \varphi(t) = \sum_{i \ge 0} c_it^i$ by the transformation $\sigma_l(t,x)=(t,t^{l}\,x)$. 
Let us note that $l$ could be chosen as positive number, because otherwise the given Laurent series $\varphi$ is already a formal power series. 
In order to keep the original and transformed system distinct, we will write $z=t^{l}\,x$. 
For the derivatives we obtain that
\[
z' = l t^{l-1} \cdot x + t^{l} \cdot x', \, \text{ etc. } 
\]
Let us define $\sigma_l(t,x_1,\ldots,x_n)=(t,t^{l}\,x_1,\ldots,t^{l}\,x_n)$. 
In $G$, we thus replace $x_j$ by $z_j:=t^{l}\,x_j$ and the derivatives by $x_j'=\frac{z_j'-l t^{l-1}z_j}{t^{l}}$, and so on. 
By taking the numerators, we obtain a system in $K[\![t]\!]\{z_1,\ldots,z_n\}$, denoted by $G^{(l)}$. 
Moreover, a formal Laurent series $\varphi=(\varphi_1,\ldots,\varphi_n)$, where all $\varphi_j$ have order greater or equal to $l$, is a solution of $G$ if and only if $(\psi_1,\ldots,\psi_n) := (t^{l}\,\varphi_1,\ldots,t^{l}\,\varphi_n)$ is a formal power series solution of $G^{(l)}$.

\subsubsection*{Puiseux series}\label{sec-Puiseux}
Let $\varphi = \sum_{i \ge 0} c_it^{(i-l)/k}$ be a Puiseux series where $c_i \in K$, $l \in \Z$, and $k>0$ is the ramification index, i.e., the minimal number such that $\varphi(t^k)$ is a formal Laurent series. 
Set $\sigma_{k,l}(t,x) = (t^{k},t^{l}\,x)$. Then $\psi:= t^{l}\,\varphi(t^k)$ is a formal power series in $t$ and we transform the derivatives as $x' = kt^{k-1+l} \cdot z' + l t^{l-1} \cdot z$, etc. 
In $G$, we thus replace $t$ by $t^k$, $x_j$ by $t^{l} \cdot z_j$ and the derivatives $x_j' = \frac{z_j'-l t^{l -1}z_j}{kt^{k-1+l}}$, etc. 
After considering just the numerators, we obtain a system $G^{(k,l)} \in R_{1,n}$ such that $(\varphi_1,\ldots,\varphi_n)$ is a formal Puiseux series solution for $G$ if and only if $(\psi_1,\ldots,\psi_n) := (t^{l}\,\varphi_1(t^k),\ldots,t^{l}\,\varphi_n(t^k))$ is a formal power series solution for $G^{(k,l)}$.

\subsubsection*{Fundamental theorem}
Let us now connect the above observations to the fundamental theorem. 
We will work with radical differential ideals $G$ because the solution set of any differential ideal and its radical coincide. For a given set of differential polynomials $\mathcal{F}$, we denote by $[\mathcal{F}]$ the smallest radical differential ideal containing $\mathcal{F}$. 
According to Ritt-Raudenbush theorem~\cite{boulier2019ritt}, every radical differential ideal is finitely generated. 
Let us now show that the generators of a radical differential ideal transform properly under the above transformation.

\begin{lem}\label{lem:generatorsaftertransformation}
Let $G$ be the radical differential ideal generated by $F_1,\ldots,F_N \in R_{1,n}$, i.e. $G=[F_1,\ldots,F_N]$. 
Then $G^{(k,l)} = [F_1^{(k,l)},\ldots,F_N^{(k,l)}]$.
\end{lem}
\begin{proof}
For the proof, we use characteristic sets representing the prime components of a given radical differential ideal~\cite{kolchin1973differential}. 
Since the order, leaders, initials and separants of $G$ and $G^{(k,l)}$ coincide (up to multiplication with monomials in $K[t]$ for the latter two), the characteristic sets of the prime components of $G$ and $G^{(k,l)}$, denoted by $\Sigma$ and $\Sigma^{(k,l)}$ respectively, are transformed into each other.
\end{proof}

In the following, let us denote by $K(\!(t)\!)= K[t^{-1}][\![t]\!]$ and $K\{\! \{t\}\!\}=\bigcup_{k \in \Z_{>0}} K(\!(t^{1/k})\!)$ the field of formal Laurent series and the field of formal Puiseux series, respectively. 
Since the transformations above change the support of the solutions, namely by a shift $l$ or a multiplication with $\tfrac{1}{k}$, we also have to include this in the solutions of the tropicalized differential systems.
Then the transformations above can be written as follows.

\begin{prop}\label{prop:solutionTransformation}
Let $K$ be a field of characteristic zero, equipped with a valuation $v_K \co K \rightarrow \T$ and $G \subset R_{1,n}$ be a differential ideal. With notations as in diagram \ref{diagram:diff-enh} for Puiseaux series over $K$, the following holds:
\[ \trop_{\widetilde v}(\Sol_{K\{\! \{t\}\!\}}(G)) = \bigcup_{l, k \in \N} \{t^{-l}A(t^k) \mid A(t) \in \trop_{\widetilde v}(\Sol_{\pk}(G^{(k,l)}))\}.
\]
\end{prop}
\begin{proof}
From the definition of $G^{(k,l)}$, as we have seen above, it follows that every formal Puiseux series solution of $G$ transforms into a formal power series solution of $G^{(k,l)}$ for some $k,l \in \Z_{>0}$.

For the converse direction, let $\psi=(\psi_1,\ldots,\psi_n)$ be a formal power series solution for some $G^{(k,l)}$ with $k,l \in \Z_{>0}$. 
Then the formal inverse $\sigma_{k,l}^{-1}(t,z):=(t^{1/k},z/t^{l})$ transforms every equation of $G^{(k,l)}$ back to an equation in $G$, where we might have to multiply with elements in $K[\![t]\!]$.
Since $\sigma_{k,l}^{-1}(t,\psi)$ defines a formal Puiseux series solution $\varphi$ of the transformed equations and, by Lemma~\ref{lem:generatorsaftertransformation}, they generate $G$, $\varphi$ is a formal Puiseux series solution of $G$.
\end{proof}

In Proposition~\ref{prop:solutionTransformation}, for the right hand side, the fundamental theorem for formal power series, Theorem \ref{thm:ft_powerseries} and \ref{thm:ft_powerseries_with_coeffs} can be applied so that we obtain the following generalization:

\begin{teorema}\label{thm:ft_PuiseuxseriesOrdinary}
Let $K$ be an uncountable, algebraically closed field of characteristic zero, equipped with a valuation $v_K \co K \rightarrow \T$. Let $G \subset R_{1,n}$ be a differential ideal. Then 
\[ 
\trop_{\widetilde v}(\Sol_{K\{\!\{t\}\!\}}(G)) = 
\bigcup_{l, k \in \N} \{t^{-l}A(t^k) \mid A(t) \in \Sol_{\T[\![t]\!]}(\trop(G^{(k,l)}))\}.
\]
\end{teorema}

\subsection{Partial case}
Let us now generalize the previous results to the multivariate case of $R_{m,n}$ with $m>1$, when the field of coefficients $K$ is equipped with the trivial valuation. 
For this purpose, let us introduce the field of formal Puiseux series in several variables as follows.

A {\em convex rational polyhedral cone} is a subset of $\R^m$ of the form
\[C = \{ \lambda_1v_1+\cdots+\lambda_rv_r \mid \lambda_i \in \R, \lambda_i \ge 0\}, \]
where $v_1,\ldots,v_r \in \Q^m$ are vectors. 
A cone is said to be {\em strongly convex} if it contains no nontrivial linear subspaces.
Let $C$ be a strongly convex rational polyhedral cone and let $\bold{d}=(d_1,\ldots,d_m) \in \Z_{>0}^m$. Set $C_\bold{d} = C \cap (\frac{1}{d_1} \cdot \Z,\ldots,\frac{1}{d_m} \cdot \Z)$. 
Using multi-index notation, the set of formal sums $\sum_{\mu \in C_\bold{d}} a_\mu t^\mu$, where $a_\mu \in K$, forms a ring and is denoted by $K[\![C_\bold{k}]\!]$. 
Let $\bold{w} \in \R_{>0}^m$ have rationally independent entries over $\Q$. 
Let $H_\bold{w}$ be the half-space $\{\bold{t} \in \R^m \mid \bold{w} \cdot \bold{t} \ge 0\}$. 
The ring of power series with support in a convex rational polyhedral cone in $H_\bold{w}$ and fractional exponents with denominators $\bold{d} \in \Z_{>0}^m$ is defined by
\[ K[\![\bold{t}]\!]_\bold{d}^\bold{w} = \bigcup_{C \subseteq H_\bold{w}} K[\![C_\bold{d}]\!]. \]
By allowing the support in some translate of $C$ we obtain the field of {\em (multivariate-) formal Laurent series}
\[K(\!(\bold{t})\!)_\bold{d}^\bold{w}= \bigcup_{\bold{l} \in (\frac{1}{d_1} \cdot \Z, \ldots, \frac{1}{d_m} \cdot \Z)} \bold{t}^{\bold{l}}\cdot K[\![\bold{t}]\!]^\bold{w}\]
and the field of {\em (multivariate) formal Puiseux series}
\[ K(\!(\bold{t})\!)_*^\bold{w} = \bigcup_{\bold{d} \in \Z_{>0}^m} K(\!(\bold{t})\!)_\bold{d}^\bold{w}.\]
Let us note that a different choice of $\bold{w}$ leads to a different representation of the Puiseux series. For a given system of differential equations, their Puiseux series solutions can thus be translated one-to-one for different $\bold{w}$'s.

Following~\cite[Section 8]{aroca2003power}, a formal Puiseux series $\varphi$ with support in a cone $\bold{l}+C_\bold{d}$ can be transformed into a formal power series with exponents in the first quadrant. 
More precisely, there exists a rational map $\kappa(\bold{t})$ with $\kappa(0)=0$, $\varphi(\kappa(\bold{t})) \in K[\![\bold{t}]\!]$ and formally invertible as a vector of Laurent monomials with differentiable inverse, found as a finite composition of ``combinatorial blow-ups'', i.e. transformations 
\[
\sigma_{I,j}(t_1,\ldots,t_m)=(s_1,\ldots,s_m) ~\text{ with }~ s_i = \begin{cases} t_i\,t_j  & i \in I, i \ne j \\ t_i & \textit{otherwise} \end{cases}
\]
where $I \subset \{1,\ldots,m\}$ contains at least two elements and $j \in I$. 
Thus, eventually we have that 
\[
\kappa(\bold{t})=(t_1 \cdot \bold{t}^{\alpha_1},\ldots,t_m \cdot \bold{t}^{\alpha_m})
\]
for some $\alpha_i \in \Z_{\ge 0}^m$ where we use the notation $\bold{t}^{\beta} = t_1^{\beta_1} \cdots t_m^{\beta_m}$. Let us denote by $\mathcal{B}$ the set of all such maps $\kappa$ that are a finite composition of combinatorial blow-ups. 
Additionally, we perform the shift by the translation vector $\bold{l}$ such that we obtain the transformation $\sigma_{\bold{d},\bold{l}}(\bold{t},\bold{x})=(\kappa(\bold{t}),\bold{t}^{\bold{l}}\,\bold{x})=:(\bold{s},\bold{z})$ for $\kappa \in \mathcal{B}, \bold{l} \in (\frac{1}{d_1} \cdot \Z, \ldots, \frac{1}{d_m} \cdot \Z)$ where $\kappa$ is depending on $\bold{w}$ and $\bold{d}$. 

Notice that the maps $\sigma_{\bold{d},\bold{l}}(\bold{t},\bold{x})$ and therefore $\kappa$ can be defined analogously on $\pb$ as all the operations in their definition involve transformations of the monomials of the power series. Also in this case we will denote them with the same symbols. 
Thus, we obtain
\[
\supp(\varphi_i(\kappa(\bold{t}))) =  \kappa(\supp(\varphi_i)).
\]

Let $G$ be a differential ideal. 
Similarly to the ordinary case, we apply the transformation $\sigma_{\bold{k},\bold{l}}(\bold{t},\bold{x})$ to $\bold{t}$, $\bold{x}$ and the partial derivatives of the components of $\bold{x}$ accordingly, namely by $\bold{x}' = \bold{t}^{-l-1}\,\kappa'^{-1} \cdot (\bold{t}\,\bold{z}'+l\,\bold{z})$ where the derivatives denote the Jacobi matrices w.r.t. $\bold{t}$. 
Let us note that $\kappa'$ is invertible and $\kappa'$ and $\kappa'^{-1}$ contain only monomials in $\Z[\bold{t}]$. 
Then a new system in $\bold{t}$ and $\bold{z}$ is obtained and we denote it, after cancelling $\bold{t}^{l}$ and multiplying with the common denominator in $K[\![\bold{t}]\!]$, by $G^{(\bold{k},\bold{l})}$.

For applying the fundamental theorem, we have to actually compute the formal inverse of the transformation $\sigma_{\bold{k},\bold{l}}$. 
The transformations $\sigma_{I,j}$ have the formal inverse $$\sigma_{I,j}^{-1}(\bold{s})= \bold{t} ~\text{ with }~ t_i= \begin{cases} s_i/s_j , & i \in I, i \ne j \\ s_i \end{cases}.$$
Thus, we obtain that $\kappa^{-1}$ is the finite decomposition of the $\sigma_{I,j}^{-1}$ and of the form $$\kappa^{-1}(\bold{s})=(s_1 \cdot \bold{s}^{\beta_1},\ldots,s_m \cdot \bold{s}^{\beta_m})$$ for some $\beta_1,\ldots,\beta_m \in \Z^m$ and $\sigma_{\bold{k},\bold{l}}^{-1}(\bold{s},\bold{z}) = (\kappa^{-1}(\bold{s}),\bold{s}^{-l} \cdot \bold{z})$. 
Since $\kappa$ is invertible such that the inverse again commutes with taking supports, we obtain
\begin{equation}\label{eq-sigmainverse}
\trop_{\widetilde v}(\varphi_i)= \kappa^{-1}(\trop_{\widetilde v}(\varphi_i(\kappa(\bold{t})))).
\end{equation}

Lemma~\ref{lem:generatorsaftertransformation} and Proposition~\ref{prop:solutionTransformation} can be generalized to the multivariate case by simply replacing $R_{1,n}$ with $R_{m,n}$ and every $k,l$ with $\bold{d},\bold{l}$ in the statements and its proofs.
Then, using the above transformation, in the case of trivial valuation, the fundamental theorem~\ref{thm:ft_PuiseuxseriesOrdinary} generalizes to the multi-variate case in the following sense.
\begin{teorema}\label{thm:ft_PuiseuxseriesPartial}
Let $K$ be an uncountable, algebraically closed field of characteristic zero. Let $G \subset R_{m,n}$ be a differential ideal and let $\bold{w} \in \R_{>0}^m$ have rationally independent entries over $\Q$. Then 

\[ 
\trop_{\widetilde v}(\Sol_{K(\!(\bold{t})\!)_*^\bold{w}}(G)) = 
\bigcup_{\bold{d},\bold{l}} \{ \sigma_{\bold{d},\bold{l}}^{-1}(A) \mid A(\bold{t}) \in \Sol_{\T[\![\bold{t}]\!]}(\trop(G^{(\bold{d},l)}))\}.
\]
where the union is taken over all $\bold{d} \in \Z_{>0}^m$ and $\bold{l} \in (\frac{1}{d_1} \cdot \Z, \ldots, \frac{1}{d_m} \cdot \Z)$.
\end{teorema}

\begin{ex}
Let us consider the differential ideal generated by the differential polynomials
\[ 
F_1 = t_1\,\frac{\partial\,x}{\partial t_1}+t_2\,\frac{\partial\,x}{\partial t_2}-x/2, \, \, \,  F_2 = \frac{\partial\,x}{\partial t_1}-\frac{\partial\,x}{\partial t_2}
\]
inside $\C[\![t_1, t_2]\!]\{x\}$.
By tropicalizing
\[\frac{\partial^{i+j}}{\partial t_1^i \partial t_2^j}F_1 = t_1\,\frac{\partial^{i+j+1}\,x}{\partial t_1^{i+1} \partial t_2^{j}}+t_2\,\frac{\partial^{i+j+1}\,x}{\partial t_1^{i} \partial t_2^{j+1}} + (i+j-1/2)\frac{\partial^{i+j}\,x}{\partial t_1^{i} \partial t_2^{j}} , \]
we directly see that $\varphi=0$ is the only possible formal power series solution. 
Let us consider for $\bold{d}=(1,1/2), \bold{l}=(0,-1/2)$ (and $\kappa(t_1,t_2)=(t_1t_2,t_2)$) the transformation $\sigma_{\bold{d},\bold{l}}(\bold{t},x)=((t_1t_2,t_2),t_2^{-1/2} \cdot x)$ such that
\[ 
\frac{\partial\,x}{\partial t_1}= \frac{1}{\sqrt{t_2}}\,\frac{\partial\,z}{\partial t_1}, \, \frac{\partial\,x}{\partial t_2} = \sqrt{t_2}\,\frac{\partial\,z}{\partial t_2}-\frac{t_1}{\sqrt{t_2}}\,\frac{\partial\,z}{\partial t_2} + \frac{1}{2\sqrt{t_2}}\,z 
\]
leads to the differential system generated by
\[ 
F_1^{(\bold{k},\bold{l})} = \frac{\partial\,z}{\partial t_2}, \, \, \, F_2^{(\bold{k},\bold{l})} = (1+t_1)\,\frac{\partial\,z}{\partial t_1}-t_2\,\frac{\partial\,z}{\partial t_2} -z/2. 
\]
For tropical solutions $A \in \B[\![t_1, t_2]\!]$ of $\trop_v(F_1^{(\bold{k},\bold{l})})$ we see that $A$ has to belong to $\B[\![t_1]\!]$. 
Similarly, by considering $\trop_v(F_2^{(\bold{k},\bold{l})})$ and all its derivatives with respect to $t_1$, a tropical solution for this system has to belong to $\B[\![t_1]\!]$. 
Indeed, the formal Puiseux series solution $\varphi = c \cdot \sqrt{t_1+t_2} = \sum_{i \ge 0} \binom{1/2}{i} \cdot t_1^it_2^{1/2-i}$ of $[F_1,F_2]$ transforms to the formal power series solution $\psi=\sum_{i \ge 0} \binom{1/2}{i} \cdot t_1^i$ of $[F_1^{(\bold{k},\bold{l})},F_2^{(\bold{k},\bold{l})}]$ which has exactly the prescribed support set, and
\[\trop_{\widetilde v}(\varphi) = \sigma_{\bold{d},\bold{l}}^{-1}(\trop_{\widetilde v}(\psi)) = \left\{ \begin{pmatrix} -1/2 \\ 1/2 \end{pmatrix}, \begin{pmatrix} 1/2 \\ -1/2 \end{pmatrix}, \ldots \right\}\]
is a solution of $\trop([F_1,F_2])$.
\end{ex}

\section{Can we generalize the fundamental theorem to series involving logarithms?}\label{sec-generalizationLogarithms}
Let us now consider series which might involve logarithms. 
For this purpose we will use (grid-based) transseries, which generalize formal Puiseux series (over $\R$). 
Informally speaking, transseries are formal Hahn series of real powers of the indeterminate $t$, exponentials, logarithms and their compositions, with real coefficients. 
It is required that the numbers of iterations of occurring exponentials and logarithms is finite and that the series are well-based. 
For a rigorous definition and details on transseries see~\cite{edgar2009transseries} or~\cite{van2006transseries}. 
A generalization to series with complex coefficients is presented in~\cite{van2001complex}. 
These series generalize formal Puiseux series over $\C$.

Let $(\mathcal{M}, \cdot)$ denote a totally ordered monomial group with powers in a real trigonometric field $K$ such as $\R$ (or $K$ to be the complexification of a real trigonometric function field such as $\C$). 
We will use $\mathcal{M}$ as the product and composition of the variables $\bold{t}$, logarithms and exponentials in $\bold{t}$. 
A grid-based transseries is then a series $\varphi = \sum_{m \in \mathcal{M}} \varphi_m \cdot m$ such that the support $\supp(\varphi)=\{m \in \mathcal{M} \mid \varphi_m \ne 0 \}$ is finitely generated in $\mathcal{M}$. We will denote as $K[\![\mathcal{M}]\!]$ the field of grid-based transseries over $\mathcal M$ with coefficients in $K$.

Let $d \co \mathcal M \rightarrow K[\![\mathcal{M}]\!]$ be a map satisfying the Leibniz rule and such that $d(\log(m)) = d(m)/m$ for all $m \in \mathcal M$. Then, by \cite[Theorem 5.1]{van2006transseries}, the map $d$ extends to an exp-log derivation on $K[\![\mathcal{M}]\!]$, thus making $(K[\![\mathcal{M}]\!],d)$ into a differential field.

Denoting with $\le$ the total order on $\mathcal M$, we can regard $\mathcal{M}\cup \{\infty\}$ as a semifield: we endow $\mathcal M$ with addition given as $\oplus := \min_{\le}$ and with multiplication given by the group operation of $\mathcal M$. These two operations are extended to $\mathcal{M}\cup \{\infty\}$ by $m \oplus \infty =m$ and $m \cdot \infty = \infty$ for all $m \in \mathcal M$. We equip $K[\![\mathcal{M}]\!]$ with a valuation $v \co K[\![\mathcal{M}]\!] \rightarrow \mathcal M$ in a tautological way by sending a power series $\varphi$ as above to the minimum of its support:
\[
    v(\varphi) := \min_\le \supp(\varphi).
\]
It is straightforward to check that the map $v$ is indeed a valuation.

Considering the idempotent semiring $\B[\![\mathcal{M}]\!]$ of Boolean transseries, we can make it into a differential semiring with a differential $d$ defined analogously as above. The map $\Phi \co (\B[\![\mathcal{M}]\!],d) \rightarrow \mathcal M$ sending a boolean transseries to the  minimum of its support in $\mathcal M$ is a tropical pair $\mathbf S$, in the language of \cite{gianmereta}. It is non-reduced (indeed its restriction to $\B(\!(t)\!)$ inside $\B[\![\mathcal{M}]\!]$ is the non-reduced tropical pair $\B(\!(t)\!) \rightarrow \Z \cup \{\infty\}$).  

Finally, we can define a map $\widetilde v \co K[\![\mathcal{M}]\!] \rightarrow \B[\![\mathcal{M}]\!]$ by coefficientwise application of the trivial valuation on the elements of $K[\![\mathcal{M}]\!]$, i.e. sending a transseries with coefficients in $K$ to its support. By definition of this map and of the differentials of its domain and codomain, it is clear that it commutes with the differentials. 
In conclusion, we built the following commutative diagram:
\[
	\begin{tikzcd}
		& \B[\![\mathcal{M}]\!]\arrow[d, "\Phi"]  \\
		 K[\![\mathcal{M}]\!] \arrow[ur, "\widetilde v"] \arrow[r, swap, "v"]& \mathcal M 
	\end{tikzcd}
\]
which satisfies all the hypothesis to be a differential enhancement of the valuation $v$, but the hypothesis of $\Phi$ being reduced. Let us denote as $\mathbf{v}=(v,\widetilde{v})\co (K[\![\mathcal{M}]\!],d) \to \mathbf{S}$ the differential enhancement above. Given a differential ideal $I \subset \diff{K[\![\mathcal{M}]\!]}{n}$, with the same proof as in \cite[Proposition 5.2.2]{gianmereta}, even if the pair $\mathbf S$ is non-reduced, we have that the following inclusion holds:
\[
\trop_{\widetilde{v}}(\textup{Sol}_{K[\![\mathcal{M}]\!]}(I)) \subseteq \textup{Sol}_\mathbf{S}(\trop_v(I)).
\]

Now, let us give an example in which the equality of the two sets above does not hold. Equivalently, the fundamental theorem does not hold for the differential enhancement $\mathbf v$ introduced above.
\begin{ex}\label{ex:fail-fund-theorem}
Consider the differential polynomial $F = tx' - x - t \in \C[\![t]\!]\{x\}$. The differential ideal generated by $F$ is prime. The solutions to the algebraic differential equation $F =0$, and thus to $[F]$, are of the form $\varphi = ct + t \log t \in \C[\![t,\log(t)]\!]$, where $c \in \C$. 
Thus, 
\[
    \trop_{\widetilde v}(\Sol(F)) = \{ t + t \log t,  t \log t \} .
\]
We can see the differential ring $\C[\![t,\log(t)]\!]$ as a differential ring $\C[\![\mathcal M]\!]$ for the totally ordered group generated by the symbols $t$ and $\log(t)$, with order given by $t \ge \log(t)$. The derivation sends $\log(t)$ to $t^{-1}$. As a semiring, $\mathcal M \cup \{\infty \}$ is isomorphic to $\Z^2 \cup \{\infty\} \subset \T_2$, thus we can write an element $t^n \log(t)^m \in \mathcal M$ as $(n,m)$ and we consider $\T_2$ to be the target of the valuation $v$.

Let us now compute the tropical differential polynomials $\trop(\{F\})$ and their solutions. 
From $\trop_v(dF) = (1,0)x'' + (0,0)$ we know that there is no solution of $\trop_v(\{F\})$ in $\B[\![t]\!]$. Since the fundamental theorem applies to solutions in $\C[\![t]\!]$, this is also telling us that there are no formal power series solutions of $[F]$.

Looking for solutions $A \in \B[\![\mathcal M]\!]$ to $ \trop_v(F) = (1,0)x' + (0,0)x + (1,0) $ we get that a solution is either of the form $A = t \log(t) + \dotsb$ or of the form $A = t  + t \log(t) + \dotsb$. 
Taking further derivatives of $F$, we obtain 
\[
\trop_v(d^k F) = (1,0)x^{(k+1)} + (0,0) x^{(k)}
\]
and evaluating $\trop_v(d^k F)$ in $A$, we obtain no further condition on the tropical solutions. We can now prove that the fundamental theorem does not hold in this context. More precisely, that given a classical solution $\varphi$, any boolean transseries of the form $A=\trop_{\widetilde v}(\varphi)+B$ with $B \in \B[\![\mathcal M]\!]$, $\Phi(B) > t \log t$ is in $\Sol(\trop_v([F]))$. 

Let us prove that $A:=\trop_{\widetilde v}(\varphi)+B$ is a solution for $\trop_v([F])$. Indeed, for every $k \in \N$, the following equality holds:
\[
    \Phi(d^k A) = \Phi(d^k \trop_{\widetilde v}(\varphi))
\]
and, as $\trop_{\widetilde v}(\varphi)$ is in $\Sol(\trop_v([F]))$, by definition of a tropical solution, every boolean transseries of the form $A=\trop_{\widetilde v}(\varphi)+B$ is in $\Sol(\trop_v([F]))$ as well.
\end{ex}

\begin{oss}
Notice that, ultimately, the fundamental theorem does not hold in this context (and analogously it does not hold when looking for solutions in $\B(\!(t)\!)$) as an essential feature of the derivation over power series is that 0 is a sink. Conversely, when looking at elements like $t^{-1}$ there is no $k \in \N$ such that $d^kt^{-1} = 0$. Even the non-reducedeness of the pair $\Phi$ does not play a role, as the same problem would arise by considering its reduction.
\end{oss}

\subsubsection*{Transformations}
Let us explain why no general transformation from transseries to formal power series, mapping the given system of algebraic differential equations to another system of algebraic differential equations fulfilling (almost all of) our assumptions in the introduction, exists. 
For this purpose, let us consider the simple case where transseries solutions $\varphi = \sum_{i \ge 0} c_i(\log(t))t^i \in K[\log(t)][\![t]\!]$ of $G \subset R_{1,n}$ are sought. 
Let us notice that we could also study $K[t][\![\exp(t)]\!]$ instead because every transseries involving only nested exponentials and logarithms can be written as a logarithm-free transseries by simply using the transformation $s(t)=\exp^{h}(t)$ where $\log^h(t)$ is the logarithmic depth.

First, $\sigma$ cannot be chosen to be rational because otherwise expressions such as $\varphi=\log(t)$ cannot be transformed into a formal power series. 
We might allow more general type of transformations as rational functions, but want to keep the other hypothesis (1),\,(2),\,(3) from the introduction. 
Then the natural choice for $\varphi=\log(t)$, namely $s(t)=\exp(t)$, is excluded by the condition $s(0)=0$. 
Also any other formal power series does not work, because for $s(t)=t \cdot s_0(t)$ with $s_0(t) \in K[\![t]\!]$, we obtain $\psi = \log(s(t)) = \log(t)+\log(s_0(t))$ which is not a formal power series since $\psi(0)$ is undefined. 
Dropping the assumption that $s(0)=0$ is problematic, because then $\varphi(s(t))$ might not be evaluated at $t=0$. 
We leave it as an open question to find $z(t,x)$ such that this issue is solved.

Second, when we use $s(t)=t$ and consider just transformations in $x$, for every constant $\varphi = c$ the evaluation $z(t,\varphi)$ has to be defined. 
If we allow more general transformations than rational functions, the natural choice  would be $z(t,x)=\exp(x/t^{k})$ for $k \in \Z$. 
In this way, $\varphi= t^k \log(t)$ would be mapped to a formal power series. 
Then, however, for $\varphi=\log(t)(t+1)$ we obtain that $z(t,\varphi) = t^{t+1}$ (for $k=0$) or $z(t,\varphi) = t^{t-1}$ (for $k=1$) which are not formal power series. 
We believe that there will not be any transformation $z$ mapping all transseries $K[\log(t)][\![t]\!]$ to formal power series.

\section{Other applications}
The transformations used for formal Laurent series and formal Puiseux series, respectively, can also be applied in other settings than that of tropical differential algebra. 
One other application, which relates to the fundamental theorem of tropical differential algebra, is given in approximation theory leading to the following result (cf.~\cite[Theorem 3.1]{denef1984power}).

\begin{teorema}\label{thm-DenefLipshitz}
Let $K \in \{ \C,\R,\Q_p \}$. Let $G$ be a finite differential system in $\Q[t]\{x_1,\ldots,x_n\}$ and let $l \in \Z$ and $ k \in \Z_{>0}$. 
Then there is an algorithm for deciding whether there is a solution $(\varphi_1,\ldots,\varphi_n) \in K(\!(t^{1/k})\!)^n$ of $G$ with $\ord(\varphi_i) \ge l$.
\end{teorema}
\begin{proof}
Let $G^{(k,l)} \in K[t]\{z_1,\ldots,z_n\}$ be the system obtained after the transformation as in Section~\ref{sec-Puiseux}. 
Solutions $(\varphi_1,\ldots,\varphi_n) \in K(\!(t^{1/k})\!)^n$ of $G$ are in one-to-one correspondence to solutions $(\psi_1,\ldots,\psi_n) \in K[\![t]\!]^n$ of $G^{(k,l)}$. 
By~\cite[Theorem 3.1]{denef1984power}, there is an algorithm for deciding the solvability of $G^{(k,l)}$ in $K[\![t]\!]$.
\end{proof}

\begin{ex}
Let us consider $F=xx'-1$ together with the initial condition $x(0)=0$.\footnote{In Theorem~\ref{thm-DenefLipshitz}, one can additionally impose a finite number of equations and inequations for the Taylor coefficients of solutions as it can also be seen in the proof of~\cite[Theorem 3.1]{denef1984power}.}
The solutions are $\varphi = \pm \sqrt{2t} \in \Q[\![t^{1/2}]\!]$, which are not formal power series. 
Let us consider $F^{(2,1)}=zz'-2t$ and $z(0)=0$. 
We see that $\psi = \pm \sqrt{2} \cdot t \in \Q[\![t]\!]$. 
This solution is easily found by~\cite[Theorem 3.1]{denef1984power}.
\end{ex}

Other results on formal power series solutions might be generalized in a similar way. 
For example, let $G$ be a given differential system and let $G^{(k,l)}$ be the transformed system as in Section~\ref{sec-Puiseux}. 
Then results on jet spaces (see e.g.~\cite{moosa2010jet}), constructed from $G^{(k,l)}$ and in one-to-one correspondence to its formal power series solutions~\cite[Remark 2.3.13]{seiler2010involution}), can be applied. 
In this way, a rigorous approximative description of the formal Puiseux series solutions of $G$ can be given. 
Moreover, following~\cite[Section 9]{seiler2010involution}, sufficiently often differentiable solutions, whose asymptotic behaviour are described by a Puiseux polynomial (they are called \textit{geometric solutions} in the reference), can be studied in this way allowing a much bigger class of solutions under investigation.


\subsection*{Acknowledgments}
First author is partially supported by the grant PID2020-113192GB-I00 (Mathematical Visualization: Foundations, Algorithms and Applications) from the Spanish MICINN and the OeAD project FR 09/2022.


\bibliographystyle{alpha}

\newcommand{\etalchar}[1]{$^{#1}$}

\end{document}